Nico is saving money for his college education. He invests some money at 7​% ,and ​$1800 less than that amount at 4%.The investments produced a total of ​$236 interest in 1 yr. How much did he invest at each​ rate?
Answer by jorel1380(3719)

Let n be the amount Nico is investing at 7%. Then, the amount at 4% would be n-1800. So:
.07n+.04(n-1800)=236
.11n-72=236
.11n=308
n=2800
n-1800=1000
☺☺☺☺